What is the centripetal acceleration of a point on the trim of flywheel 1.20 m in diameter, turning at a rate of 1,200 r/min?


1
Expert's answer
2021-09-20T09:59:41-0400

The centripetal acceleration of a point on a rotating object is given as follows:


"a= \dfrac{v^2}{R}"

where "v" is the linear speed of the point and "R = 1.20m\/2=0.6m" is the distance from the center of rotation to the point.

Let the frequency be "f = 1200r\/min = 20Hz". Then the speed is:


"v = 2\pi fR"

Thus, obtain:


"a = \dfrac{(2\pi fR)^2}{R} = 4\pi^2 f^2R\\
a = 4\pi^2\cdot (20Hz)^2\cdot 0.6m \approx 9.47\times 10^3m/s^2"

Answer. "9.47\times 10^3m/s^2".